### v2

# Harmonic higher and extended weight enumerators

Abstract.
In this paper, we present the harmonic generalizations of well-known polynomials of codes over finite fields, namely the higher weight enumerators and the extended weight enumerators, and we derive the correspondences between these weight enumerators. Moreover, we present the harmonic generalization of Greene’s Theorem for the higher (resp. extended) weight enumerators. As an application of this Greene’s-type theorem, we provide the MacWilliams-type identity for harmonic higher weight enumerators of codes over finite fields. Finally, we use this new identity to give a new proof of the Assmus-Mattson Theorem for subcode supports of linear codes over finite fields using harmonic higher weight enumerators.
*Corresponding author
Key Words: Tutte polynomials, weight enumerators, matroids, codes, harmonic functions.
2010 Mathematics Subject Classification. Primary 11T71; Secondary 94B05, 11F11.
1. Introduction
The classification of codes is one of the most challenging problems in the theory of algebraic coding for the last few decades. Mathematicians found a way to solve this problem when MacWilliams [20] introduced a powerful identity for codes in 1963, now known as the MacWilliams identity. This identity plays an important role in classifying codes. The identity also allowed Assmus and Mattson [1] to prove a famous theorem now known as the Assmus-Mattson Theorem that describes how linear codes can generate $t$-designs.
In the decade prior to these results, Tutte [25] presented a two-variable polynomial of a graph, since known as the Tutte polynomial, that not only can reveal the internal structure of a graph but also can be used to determine several other graph polynomials; see also [26]. Later, Crapo [12] gave an appropriate generalization of the Tutte polynomial by defining it for matroids. Greene [16] presented a rare and unexpected relation between the weight enumerator of a linear code over finite field and the Tutte polynomial of a matroid. Moreover, using this relation, Greene [16] also gave a combinatorial proof of the MacWilliams identity for the weight enumerator of a code. Wei [27] introduced the notion of subcode weights in the study of coding theory, and Kløve [18] gave a remarkable generalization of the MacWilliams identity for these subcode weights. Many authors have studied Greene’s Theorem between the linear codes and matroids; for instance, see [11]. Of note, Britz [5] generalized Greene’s Theorem with respect to the rank generating function of a matroid and the higher support weight enumerator of a linear code over a finite field. By applying this research, Britz and Shiromoto [8] generalized the Assmus-Mattson Theorem with respect to subcode supports of linear codes over finite fields; see also [4, 9, 10, 24].
Delsarte [14] introduced discrete harmonic functions on a finite set. Later, Bachoc [2] obtained a striking generalization of the MacWilliams identity by introducing the concept of the harmonic weight enumerators of a binary code. The harmonic weight enumerator of code is a weight enumerator of a code associated to a discrete harmonic function. Recently, Chakraborty, Miezaki and Oura [11] introduced harmonic Tutte polynomials of matroids, and presented the harmonic generalization of Greene’s Theorem. Subsequently, a demi-matroid analogue of harmonic Tutte polynomials was discussed in [7].
In this paper, we introduce the notion of harmonic higher (resp. extended) weight enumerators of codes over finite fields. We present fundamental theorems regarding these weight enumerators, including their Greene’s-type theorem. As an application of this theorem, we provide the MacWilliams-type identity for harmonic higher weight enumerators over finite fields. Finally, we use this new identity for harmonic higher weight enumerators to give a new proof of the Assmus-Mattson Theorem for subcode supports of linear codes over finite fields.
This paper is organized as follows. In Section 2, we present basic definitions and properties that are frequently used in this paper, of discrete harmonic functions, linear codes over finite fields and matroids. In Section 3, we define the harmonic higher (resp. extended) weight enumerators for codes over finite fields, and give formulae to compute these; see Theorems 3.2 and 3.4. We also provide the relationships between these weight enumerators; see Theorems 3.5 and 3.6. Moreover, we present the MacWilliams-type identity for harmonic higher (resp. extended) weight enumerators of codes over finite fields; see Theorems 3.1 and 3.3. In Section 4, we give the Greene-type identity for the harmonic higher (resp. extended) weight enumerators for codes over finite fields; see Theorems 4.1 and 4.2. We apply the Greene-type identity for harmonic extended weight enumerators of codes over finite fields to prove the MacWilliams-type identity for the harmonic higher weight enumerators of codes over finite fields. Finally, in Section 5, we use harmonic higher weight enumerators to give a new proof of the Assmus-Mattson Theorem for subcode supports of linear codes over finite fields; see Theorem 5.2 and its proof.
2. Preliminaries
In this section, we present basic definitions and notation for linear codes and matroids that are frequently needed in this paper. We mostly follow the definitions and notation of [13, 21, 22]. We also present definitions and properties of (discrete) harmonic functions, following [2, 3, 14] .
2.1. Discrete harmonic functions
Let $E:=\{1,2,\ldots,n\}$ be a finite set. We define $E_{d}:=\{X\subseteq E:|X|=d\}$ for $d=0,1,\ldots,n$. The set of all subsets of $E$ is denoted by $2^{E}$. We denote by $\mathbb{R}2^{E}$ and $\mathbb{R}E_{d}$ the real vector spaces spanned by the elements of $2^{E}$ and $E_{d}$, respectively. An element of $\mathbb{R}E_{d}$ is denoted by
(1) $f:=\sum_{Z\in E_{d}}f(Z)Z$
and is identified with the real-valued function on $E_{d}$ given by $Z\mapsto f(Z)$. Such an element $f\in\mathbb{R}E_{d}$ can be extended to an element $\widetilde{f}\in\mathbb{R}2^{E}$ by setting, for all $X\in 2^{E}$,
(2) $\widetilde{f}(X):=\sum_{Z\in E_{d},Z\subseteq X}f(Z).$
Note that $\widetilde{f}(\emptyset)=f(\emptyset)$ when $d=0$, and that $\widetilde{f}(\emptyset)=0$ otherwise. If an element $g\in\mathbb{R}2^{E}$ is equal to $\widetilde{f}$ for some $f\in\mathbb{R}E_{d}$, then we say that $g$ has degree $d$. The differentiation operator $\gamma$ on $\mathbb{R}E_{d}$ is defined by linearity from the identity
(3) $\gamma(Z):=\sum_{Y\in{E}_{d-1},Y\subseteq Z}Y$
for all $Z\in E_{d}$ and for all $d=0,1,\ldots,n$. Also, $\operatorname{Harm}_{d}(n)$ is the kernel of $\gamma$:
(4) $\operatorname{Harm}_{d}(n):=\ker\left(\gamma\big{|}_{\mathbb{R}E_{d}}\right).$
Remark 2.1 ([2, 14]).
Let $f\in\operatorname{Harm}_{d}(n)$ and $i\in\{0,1,\ldots,d-1\}$. Then $\gamma^{d-i}(f)=0$. This means from definition (3) that
$\sum_{X\in E_{i}}\left(\sum_{\begin{subarray}{c}Z\in E_{d},X\subseteq Z\end{subarray}}f(Z)\right)X=0.$
This implies that $\sum_{\begin{subarray}{c}Z\in E_{d},X\subseteq Z\end{subarray}}f(Z)=0$ for any $X\in E_{i}$.
Remark 2.2.
Let $f\in\operatorname{Harm}_{d}(n)$. Since $\sum_{Z\in E_{d}}f(Z)=0$, it is easy to check from (3) that $\sum_{X\in E_{t}}\widetilde{f}(X)=0$, where $1\leq d\leq t\leq n$.
Example 2.1.
Let $E=\{1,2,3,4\}$ and $d=2$. Let $f\in\mathbb{R}E_{2}$ be the element
$f=a_{1}\{1,2\}+a_{2}\{1,3\}+a_{3}\{1,4\}+a_{4}\{2,3\}+a_{5}\{2,4\}+a_{6}\{3,4\}.$
If $X=\{1,3,4\}$, then $\widetilde{f}(X)=a_{2}+a_{3}+a_{6}$. Now applying differential operator $\gamma$ on $f$, we have
$\gamma(f)=(a_{1}+a_{2}+a_{3})\{1\}+(a_{1}+a_{4}+a_{5})\{2\}+(a_{2}+a_{4}+a_{6})\{3\}+(a_{3}+a_{5}+a_{6})\{4\}.$
It follows that $f\in\operatorname{Harm}_{2}(4)$ if and only if $a_{3}=a_{4}=-a_{1}-a_{2}$, $a_{5}=a_{2}$ and $a_{6}=a_{1}$.
2.2. Linear codes
Let $\mathbb{F}_{q}$ be a finite field of order $q$, where $q$ is a prime power. Then $\mathbb{F}_{q}^{n}$ denotes the vector space of dimension $n$ with the usual inner product: $\mathbf{u}\cdot\mathbf{v}:=u_{1}v_{1}+\cdots+u_{n}v_{n}$ for $\mathbf{u},\mathbf{v}\in\mathbb{F}_{q}^{n}$, where $\mathbf{u}=(u_{1},\ldots,u_{n})$ and $\mathbf{v}=(v_{1},\ldots,v_{n})$. For $\mathbf{u}\in\mathbb{F}_{q}^{n}$, we call $\operatorname{supp}(\mathbf{u}):=\{i\in E\mid u_{i}\neq 0\}$ the support of $\mathbf{u}$, and $\operatorname{wt}(\mathbf{u}):=|\operatorname{supp}(\mathbf{u})|$ the weight of $\mathbf{u}$. Similarly, the support and weight of each subset of vectors $D\subseteq\mathbb{F}_{q}^{n}$ are defined as follows:
$\operatorname{Supp}(D)$ $:=\bigcup_{\mathbf{u}\in D}\operatorname{supp}(\mathbf{u})$
$\operatorname{wt}(D)$ $:=|\operatorname{Supp}(D)|.$
An $\mathbb{F}_{q}$-linear code of length $n$ is a linear subspace of $\mathbb{F}_{q}^{n}$. Moreover, an $[n,k,d_{\min}]$ code is an $\mathbb{F}_{q}$-linear code of length $n$ with dimension $k$ and minimum weight $d_{\min}$. Frequently, we call $[n,k,d_{min}]$ codes $[n,k]$ codes. The dual code $C^{\perp}$ of an $\mathbb{F}_{q}$-linear code $C$ can be defined as follows:
$C^{\perp}:=\{\mathbf{u}\in\mathbb{F}_{q}^{n}\mid\mathbf{u}\cdot\mathbf{v}=0\mbox{ for all }\mathbf{v}\in C\}.$
Let $C$ be an $[n,k]$ code. Let $r,i$ be positive integers such that $r\leq k$ and $i\leq n$. Now we define
$\mathcal{D}_{r}(C)$ :=$:=\{D\mid D\text{ is an }[n,r]\text{ subcode of }C\},$
$\mathcal{S}_{r}(C)$ :=$:=\{\operatorname{Supp}(D)\mid D\in\mathcal{D}_{r}(C)\},$
$\mathcal{S}_{r,i}(C)$ :=$:=\{X\in\mathcal{S}_{r}(C)\mid|X|=i\}.$
Note that, in general, the sets $\mathcal{S}_{r}(C)$ and $\mathcal{S}_{r,i}(C)$ are multisets. Also, the support of a nonzero codeword $\mathbf{v}$ is identical to that of its span $\langle\mathbf{v}\rangle$. Conversely, each member of $\mathcal{S}_{1}(C)$ is the support of some nonzero codeword. Therefore, the sets $\mathcal{S}_{1}(C),\ldots,\mathcal{S}_{k}(C)$ extend the notion of codeword support.
Let $C$ be an $\mathbb{F}_{q}$-linear code of length $n$. Then the $r$-th generalized Hamming weight of $C$ for any $r$, $1\leq r\leq k$, is defined as
$d_{r}:=d_{r}(C):=\min\{\operatorname{wt}(D)\mid D\in\mathcal{D}_{r}(C)\}.$
Moreover, the $r$-th higher weight distribution of $C$ is the sequence
$\{A_{i}^{(r)}(C)\mid i=0,1,\dots,n\},$
where $A_{i}^{(r)}(C)$ is the number of subcodes with a given weight $i$ and dimension $r$. That is,
$A_{i}^{(r)}(C):=\#\{D\in\mathcal{D}_{r}(C)\mid\operatorname{wt}(D)=i\}.$
Then the polynomial
$W_{C}^{(r)}(x,y):=\sum^{n}_{i=0}A_{i}^{(r)}(C)x^{n-i}y^{i}$
is called the $r$-th higher weight enumerator of $C$ and satisfies the following MacWilliams-type identity stated in [17, Theorem 5.14]:
$W_{C^{\perp}}^{(r)}(x,y)=\sum_{j=0}^{r}\sum_{\ell=0}^{j}(-1)^{r-j}\frac{q^{{r-j\choose 2}-j(r-j)-\ell(j-\ell)-jk}}{[r-j]_{q}[j-\ell]_{q}}W_{C}^{(\ell)}(x+(q^{j}-1)y,x-y).$
Here, for all integers $a,b\geq 0$,
$[a]_{q}:=[a,a]_{q}\qquad\textrm{where}\qquad[a,b]_{q}:=\prod_{i=0}^{b-1}\big{(}q^{a}-q^{i}\big{)}.$
Let $C$ be an $\mathbb{F}_{q}$-linear code of length $n$. Then for an arbitrary subset $X\subseteq E$ and integer $r\geq 0$, we define
$C(X)$ :=$:=\{\mathbf{u}\in C\mid u_{i}=0\text{ for all }i\in X\}\,,$
$\ell(X)$ $:=\dim C(X)\,,$
$B_{X}^{(r)}(C)$ :=#$:=\#\{D\subseteq C(X)\mid D\text{ is a subspace of dimension }r\}\,.$
Also, define
${\left[\!\begin{array}[]{c}a\\ b\end{array}\!\right]}_{q}$ $:=\frac{[a,b]_{q}}{[b]_{q}}=\prod_{i=0}^{b-1}\frac{q^{a}-q^{i}}{q^{b}-q^{i}}\,.$
The following two lemmas provide important and useful identities. The second lemma follows immediately from [19, Theorem 25.2].
Lemma 2.1.
${B_{X}^{(r)}(C)=\left[\!\begin{array}[]{c}\ell(X)\\ r\end{array}\!\right]}_{q}$.
Lemma 2.2.
$[a,b]_{q}=\displaystyle\sum_{i=0}^{b}{\left[\!\begin{array}[]{c}b\\ i\end{array}\!\right]}_{q}(-1)^{b-i}q^{\binom{b-i}{2}}(q^{a})^{i}$.
2.3. Matroids
Matroids can be defined in several equivalent ways. We prefer the definition which is in terms of independent sets. A (finite) matroid $M$ is an ordered pair $(E,\mathcal{I})$ consisting of set $E$ and a collection $\mathcal{I}$ of subsets of $E$ satisfying the following conditions:
  • (I1)
    $\emptyset\in\mathcal{I}$,
  • (I2)
    if $I\in\mathcal{I}$ and $J\subset I$, then $J\in\mathcal{I}$, and
  • (I3)
    if $I,J\in\mathcal{I}$ with $|I|<|J|$, then there exists $j\in J\setminus I$ such that $I\cup\{j\}\in\mathcal{I}$.
The elements of $\mathcal{I}$ are called the independent sets of $M$, and $E$ is called the ground set of $M$. A subset of the ground set $E$ that does not belong to $\mathcal{I}$ is called dependent. An independent set is called a basis if it is not contained in any other independent set. The dual matroid of $M$ is denoted by $M^{\ast}$ and has as its independent sets all subsets of the complements of each basis of $M$.
It follows from axiom (I3) that the cardinalities of all bases in a matroid $M$ are equal; this cardinality is called the rank of $M$. The rank $\rho(X)$ of an arbitrary subset $X$ of $E$ is the size of the largest independent subset of $X$. That is, $\rho(X):=\max\{|I|:I\in\mathcal{I}\text{ and }I\subseteq X\}.$ In particular, $\rho(\emptyset)=0$. We call $\rho(E)$ the rank of $M$. We refer the readers to [22] for more information on matroids.
Definition 2.1.
Let $M$ be a matroid on the set $E$ having a rank function $\rho$. The Tutte polynomial of $M$ is defined as follows:
$T(M;x,y):={\sum_{X\subseteq E}}(x-1)^{\rho(E)-\rho(X)}(y-1)^{|X|-\rho(X)}.$
Let $A$ be a $k\times n$ matrix over a finite field $\mathbb{F}_{q}$. This gives a matroid $M[A]$ on the set $E$ in which a set $I$ is independent if and only if the family of columns of $A$ whose indices belong to $I$ is linearly independent. Such a matroid is called a vector matroid. Let $G$ be a generator matrix of an $\mathbb{F}_{q}$-linear code $C$ of length $n$. Then the vector matroid $M_{C}:=M[G]$ is a matroid on $E$. Note that $M_{C}$ independent of the choice of the generator matrix $G$. Moreover, it is well known that the dual matroid corresponds to dual code: $(M_{C})^{\ast}=M_{C^{\perp}}$.
3. Coding theory associated to harmonic functions
Bachoc [2] first defined the notion of harmonic weight enumerators for binary codes and presented a MacWilliams-type identity for these enumerators. Later, several articles extended this concept for linear codes over finite fields [3, 24] as well as over finite rings [7]. In this section, we introduce the harmonic generalization of higher (resp. extended) weight enumerators for codes over finite fields.
Definition 3.1.
Let $C$ be an $\mathbb{F}_{q}$-linear code of length $n$. Let $f\in\operatorname{Harm}_{d}(n)$, where $d\neq 0$. Then the harmonic weight enumerator of $C$ associated to $f$ is defined as follows:
$W_{C,f}(x,y):=\sum_{i=0}^{n}A_{i,f}(C)x^{n-i}y^{i},$
where
$A_{i,f}(C):=\sum_{\mathbf{u}\in C,\operatorname{wt}(\mathbf{u})=i}\widetilde{f}(\operatorname{supp}(\mathbf{u})).$
3.1. Harmonic higher weight enumerators
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$ and let $f\in\operatorname{Harm}_{d}(n)$, where $d\neq 0$.
Definition 3.2.
The harmonic $r$-th higher weight enumerator of $C$ associated to $f$ is defined as follows:
$W_{C,f}^{(r)}(x,y):=\sum^{n}_{i=0}A_{i,f}^{(r)}(C)x^{n-i}y^{i},$
where
$A_{i,f}^{(r)}(C):=\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D)=i\end{subarray}}\widetilde{f}(\operatorname{supp}(D))\,.$
Remark 3.1.
Clearly, $A_{0,f}^{(r)}(C)=0$ for all $0\leq r\leq k$.
Remark 3.2.
Since every $1$-dimensional subspace of $C$ contains $q-1$ nonzero codewords, $(q-1)A_{i,f}^{(1)}(C)=A_{i,f}(C)$ for $0\leq i\leq n$. Therefore,
$W_{C,f}(x,y)=W_{C,f}^{(1)}(x,y)\,.$
Remark 3.3.
If $\deg f=0$, then $W_{C,f}^{(r)}(x,y)=f(\emptyset)\,W_{C}^{(r)}(x,y)$.
Lemma 3.1 ([2]).
Let $f\in\operatorname{Harm}_{d}(n)$ and $J\subseteq E$, and define
${f^{(i)}(J):=\sum_{\begin{subarray}{c}Z\in E_{d},\\ |J\cap Z|=i\end{subarray}}f(Z).}$
Then for all $0\leq i\leq d$, $f^{(i)}(J)=(-1)^{d-i}\binom{d}{i}\widetilde{f}(J)$.
Remark 3.4.
From the definition of $\widetilde{f}$ for $f\in\operatorname{Harm}_{d}(n)$, we have $\widetilde{f}(J)=0$ for each $J\in 2^{E}$ such that $|J|<d$. Let $J\in 2^{E}$ and set $I=E\setminus J$. Then by Lemma 3.1,
${\widetilde{f}(J)=\sum_{\begin{subarray}{c}Z\in E_{d},\\ Z\subset J\end{subarray}}f(Z)=\sum_{\begin{subarray}{c}Z\in E_{d},\\ |I\cap Z|=0\end{subarray}}f(Z)=f^{(0)}(I)=(-1)^{d}\widetilde{f}(E\setminus J).}$
Moreover, we have from the above equality that if $|J|>n-d$, then $\widetilde{f}(J)=0$.
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$. Then it is immediate from Remark 3.4 that the harmonic $r$-th higher weight enumerator $W_{C,f}^{(r)}(x,y)$ of $C$ associated to $f\in\operatorname{Harm}_{d}(n)$ for any $r$, $1\leq r\leq k$ can be written as:
${W_{C,f}^{(r)}(x,y)=(xy)^{d}Z_{C,f}^{(r)}(x,y),}$
where $Z_{C,f}^{(r)}$ is a homogeneous polynomial of degree $n-2d$ of the form as follows:
${Z_{C,f}^{(r)}(x,y)=\sum_{i=d}^{n-d}A_{i,f}^{(r)}(C)x^{n-i-d}y^{i-d}.}$
Example 3.1.
Let $C$ be the binary linear code of length $3$ with codewords
$(0,0,0),(0,0,1),(1,1,0),(1,1,1)\,.$
We consider $f\in\operatorname{Harm}_{1}(3)$, where $f=a\{1\}+b\{2\}-(a+b)\{3\}$. Then by direct computation, we find the harmonic weight enumerator of $C$ associated to $f$ to be
$W_{C,f}^{(1)}=-(a+b)x^{2}y+(a+b)xy^{2}=(xy)^{1}Z_{C,f}^{(1)},$
where $Z_{C,f}^{(1)}=(a+b)(y-x)$.
Theorem 3.1 (MacWilliams type identity).
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$, and let $f\in\operatorname{Harm}_{d}(n)$. Then
$Z_{C^{\perp},f}^{(r)}(x,y)=\sum_{j=0}^{r}\sum_{\ell=0}^{j}(-1)^{r+d-j}$ $\frac{q^{{r-j\choose 2}-j(r-j)-\ell(j-\ell)-j(k-d)}}{[r-j]_{q}[j-\ell]_{q}}$
$Z_{C,f}^{(\ell)}\left({x+(q^{j}-1)y},{x-y}\right).$
To pave the way for the proof of Theorem 3.1 that we give in Section 4, we begin by presenting some propositions and theorems related to $Z_{C,f}^{(r)}(x,y)$.
Proposition 3.1.
Let $f\in\operatorname{Harm}_{d}(n)$ and $X\subseteq E$. Define
$B_{t,f}^{(r)}(C):=\sum_{X\in E_{t}}\widetilde{f}(X)B_{X}^{(r)}(C).$
Then we have the following relation between $B_{t,f}^{(r)}(C)$ and $A_{i,f}^{(r)}(C)$:
$B_{t,f}^{(r)}(C)=(-1)^{d}\sum_{i=d}^{n-t}{\binom{n-d-i}{t-d}}A_{i,f}^{(r)}(C)$
if $d\leq t\leq n-d$; otherwise, $B_{t,f}^{(r)}(C)=0$.
Proof.
It is immediate from (2) and Remark 3.4 that $B_{t,f}^{(r)}(C)=0$ for $0\leq t<d$ and $n-d<t\leq n$. We now focus on $t$ with $d\leq t\leq n-d$. By Remark 3.4, we have
$B_{t,f}^{(r)}(C)$ $=\sum_{X\in E_{t}}\widetilde{f}(X)B_{X}^{(r)}(C)$
$=(-1)^{d}\sum_{X\in E_{t}}\widetilde{f}(E\setminus X)B_{X}^{(r)}(C)$
$=(-1)^{d}\sum_{X\in E_{t}}\,\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{Supp}(D)\subseteq E\setminus X\end{subarray}}\widetilde{f}(E\setminus X)$
$=(-1)^{d}\sum_{D\in\mathcal{D}_{r}(C)}\,\,\sum_{\begin{subarray}{c}X\in E_{t},\\ X\subseteq E\setminus\operatorname{Supp}(D)\end{subarray}}\widetilde{f}(E\setminus X)$
$=(-1)^{d}\sum_{i=d}^{n-t}\,\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D)=i\end{subarray}}\,\,\sum_{\begin{subarray}{c}X\in E_{t},\\ X\subseteq E\setminus\operatorname{Supp}(D)\end{subarray}}\widetilde{f}(E\setminus X)$
$=(-1)^{d}\sum_{i=d}^{n-t}\,\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D)=i\end{subarray}}\,\,\sum_{\begin{subarray}{c}Y\in E_{n-i-t},\\ Y\subseteq E\setminus\operatorname{Supp}(D)\end{subarray}}\widetilde{f}(\operatorname{Supp}(D)\cup Y)$
$=(-1)^{d}\sum_{i=d}^{n-t}\,\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D)=i\end{subarray}}\,\,\sum_{\begin{subarray}{c}Y\in E_{n-i-t},\\ Y\subseteq E\setminus\operatorname{Supp}(D)\end{subarray}}\,\,\sum_{\begin{subarray}{c}Z\in E_{d},\\ Z\subseteq\operatorname{Supp}(D)\cup Y\end{subarray}}f(Z)$
$=(-1)^{d}\sum_{i=d}^{n-t}\,\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D)=i\end{subarray}}\,\,\sum_{j=0}^{d}\,\,\sum_{\begin{subarray}{c}Z\in E_{d},\\ |Z\cap\operatorname{Supp}(D)|=j\end{subarray}}\,\,\sum_{\begin{subarray}{c}W\in E_{n-i-t-d+j},\\ W\subseteq E\setminus(\operatorname{Supp}(D)\cup Z)\end{subarray}}f(Z)$
$=(-1)^{d}\sum_{i=d}^{n-t}\,\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D)=i\end{subarray}}\,\,\sum_{j=0}^{d}\binom{n-i-d+j}{n-i-t-d+j}\sum_{\begin{subarray}{c}Z\in E_{d},\\ |\operatorname{Supp}(D)\cap Z|=j\end{subarray}}f(Z)\,.$
Using inclusion-exclusion to count the $t$-element subsets of an $x$-element set that include some fixed $d$ elements gives the identity
(5) $\sum_{j=0}^{d}(-1)^{j}\binom{x-j}{t}\binom{d}{j}=\binom{x-d}{t-d}\,.$
By Lemma 3.1 and (5), it follows that
$B_{t,f}^{(r)}(C)$ $=(-1)^{d}\sum_{i=d}^{n-t}\,\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D)=i\end{subarray}}\,\sum_{j=0}^{d}\binom{n-i-d+j}{t}(-1)^{d-j}\binom{d}{j}\widetilde{f}(\operatorname{Supp}(D))$
$=(-1)^{d}\sum_{i=d}^{n-t}\,\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D)=i\end{subarray}}\,\sum_{j=0}^{d}(-1)^{j}\binom{n-i-j}{t}\binom{d}{j}\widetilde{f}(\operatorname{Supp}(D))$
$=(-1)^{d}\sum_{i=d}^{n-t}\,\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D)=i\end{subarray}}\binom{n-d-i}{t-d}\widetilde{f}(\operatorname{Supp}(D))$
$=(-1)^{d}\sum_{i=d}^{n-t}\binom{n-d-i}{t-d}A_{i,f}^{(r)}(C)\,.\qed$
Now we have the following result.
Theorem 3.2.
Let $C$ be an $[n,k]$ code $C$ over $\mathbb{F}_{q}$, and let $f\in\operatorname{Harm}_{d}(n)$. Then
$Z_{C,f}^{(r)}(x,y)=(-1)^{d}\sum_{t=d}^{n-d}B_{t,f}^{(r)}(C)(x-y)^{t-d}y^{n-t-d}\,.$
Proof.
By Proposition 3.1 and the binomial expansion of $x^{n-i}=((x-y)+y)^{n-i}$, we have
$(-1)^{d}$ $\sum_{t=d}^{n-d}B_{t,f}^{(r)}(C)(x-y)^{t-d}y^{n-t-d}$
$=$ $\sum_{t=d}^{n-d}\,\sum_{i=d}^{n-t}{\binom{n-d-i}{t-d}}A_{i,f}^{(r)}(C)(x-y)^{t-d}y^{n-t-d}$
$=$ $\sum_{i=d}^{n-d}A_{i,f}^{(r)}(C)\left(\sum_{t=d}^{n-i}{\binom{n-d-i}{t-d}}(x-y)^{t-d}y^{(n-d-i)-(t-d)}\right)y^{i-d}$
$=$ $\sum_{i=d}^{n-d}A_{i,f}^{(r)}(C)x^{n-d-i}y^{i-d}$
$=$ $\sum_{i=0}^{n}A_{i,f}^{(r)}(C)x^{n-i-d}y^{i-d}$
$=$ $\,Z_{C,f}^{(r)}(x,y),$
since $A_{i,f}^{(r)}(C)=0$ for $i<d$ and $i>n-d$. ∎
3.2. Harmonic extended weight enumerators
Let $m$ be a positive integer. Then any $[n,k]$ code $C$ over $\mathbb{F}_{q}$ can be extended to a linear code $C{[q^{m}]}:=C\otimes\mathbb{F}_{q^{m}}$ over $\mathbb{F}_{q^{m}}$ by considering all the $\mathbb{F}_{q^{m}}$-linear combinations of the codewords in $C$. We call $C{[q^{m}]}$ the $\mathbb{F}_{q^{m}}$-extension code of $C$. Clearly, $\dim_{\mathbb{F}_{q}}C(X)=\dim_{\mathbb{F}_{q^{m}}}C[q^{m}](X)$ for all $X\subseteq E$ (see [17, p. 34]).
Definition 3.3.
Let $C$ be an $[n,k]$ code $C$ over $\mathbb{F}_{q}$, and let $f\in\operatorname{Harm}_{d}(n)$. Then the harmonic $q^{m}$-extended weight enumerator of $C$ associated to $f$ can be defined as follows:
$W_{C,f}(x,y;q^{m}):=\sum_{i=0}^{n}A_{i,f}^{q^{m}}(C)x^{n-i}y^{i},$
where
$A_{i,f}^{q^{m}}(C):=\sum_{\mathbf{u}\in C{[q^{m}]},\operatorname{wt}(\mathbf{u})=i}\widetilde{f}(\operatorname{Supp}(\mathbf{u}))\,.$
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$. Then by Remark 3.4 one can easily observe that, for any positive integer $m$, the harmonic $q^{m}$-extended weight enumerator $W_{C,f}(x,y;q^{m})$ of $C$ associated to $f\in\operatorname{Harm}_{d}(n)$ can be written as
${W_{C,f}(x,y;q^{m})=(xy)^{d}Z_{C,f}(x,y;q^{m}),}$
where $Z_{C,f}(x,y;q^{m})$ is the following homogeneous polynomial of degree $n-2d$:
${Z_{C,f}(x,y;q^{m})=\sum_{i=d}^{n-d}A_{i,f}^{q^{m}}(C)x^{n-i-d}y^{i-d}.}$
Now we have the following harmonic analogue of the MacWilliams identity for $\mathbb{F}_{q^{m}}$-extended code. Bachoc [2] and Tanabe [24] presented the harmonic generalizations of the MacWilliams identity for binary and non-binary codes, respectively. Later, an alternative approach of the proof of this identity was given in [11].
Theorem 3.3 (MacWilliams type identity).
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$, and let $f\in\operatorname{Harm}_{d}(n)$. Then
$Z_{C^{\perp},f}(x,y;q^{m})=(-1)^{d}q^{-(k-d)m}Z_{C,f}\left({x+(q^{m}-1)y},{x-y};q^{m}\right).$
Following the approach given in [11], we provide the proof of Theorem 3.3 in Section 4. In order to do that we need to present some necessary propositions and theorems related to $Z_{C,f}(x,y;q^{m})$.
Definition 3.4.
Let $C$ be an $\mathbb{F}_{q}$-linear code of length $n$ and let $f\in\operatorname{Harm}_{d}(n)$. For any subset $X\subseteq E$ and positive integer $m$, we define
$B_{X}^{q^{m}}(C)$ $:=(q^{m})^{\ell(X)}-1,$
$B_{t,f}^{q^{m}}(C)$ $:=\sum_{X\in E_{t}}\widetilde{f}(X)B_{X}^{q^{m}}(C).$
Lemma 3.2.
For all $d\leq t\leq n-d$,
$\sum_{X\in E_{t}}(q^{m})^{\ell(X)}=B_{t,f}^{q^{m}}(C)\,.$
Proof.
Since $\sum_{X\in E_{t}}\widetilde{f}(X)=0$ for $d\leq t\leq n-d$, we can write
$\sum\limits_{X\in E_{t}}\widetilde{f}(X)(q^{m})^{\ell(X)}$ $={\sum\limits_{X\in E_{t}}\widetilde{f}(X)(B_{X}^{q^{m}}(C)+1)}$
$=B_{t,f}^{q^{m}}(C)+\sum\limits_{X\in E_{t}}\widetilde{f}(X)$
$=B_{t,f}^{q^{m}}(C)\,.\qed$
Proposition 3.2.
The relation between $B_{t,f}^{q^{m}}(C)$ and $A_{i,f}^{q^{m}}(C)$ is as follows:
$B_{t,f}^{q^{m}}(C)=(-1)^{d}\sum_{i=d}^{n-t}\binom{n-d-i}{t-d}A_{i,f}^{q^{m}}(C),$
if $d\leq t\leq n-d$; otherwise, $B_{t,f}^{q^{m}}(C)=0$.
The proof of the above proposition is similar to the proof of [11, Proposition 3.8] and is therefore omitted. Using the above proposition, we have the extended analogue of [11, Proposition 3.9] as follows. Since the proof is straightforward, we leave it to the reader.
Theorem 3.4.
Let $C$ be an $\mathbb{F}_{q}$-linear code of length $n$, and let $f\in\operatorname{Harm}_{d}(n)$. Then
$Z_{C,f}(x,y;q^{m})=(-1)^{d}\sum_{t=d}^{n-d}B_{t,f}^{q^{m}}(C)(x-y)^{t-d}y^{n-t-d}.$
Using Lemma 3.2, we have the following proposition.
Proposition 3.3.
We can express $Z_{C,f}(x,y;q^{m})$ as follows:
$Z_{C,f}(x,y;q^{m})=(-1)^{d}\sum\limits_{t=d}^{n-d}\,\sum\limits_{X\in E_{t}}\widetilde{f}(X)(q^{m})^{\ell(X)}(x-y)^{t-d}y^{n-t-d}.$
3.3. The link between the two weight enumerators
In this section, we establish two-way relationships between the harmonic version of two weight polynomials, namely extended weight enumerators and higher weight enumerators. Each of these relationships plays an important role in the development of the subsequent sections.
Let $\operatorname{Mat}(m\times n,\mathbb{F}_{q})$ be the $\mathbb{F}_{q}$-linear space of $m\times n$ matrices with entries in $\mathbb{F}_{q}$. Then for an $\mathbb{F}_{q}$-linear code $C$ of length $n$, $S_{m,n}(C)$ denotes the linear subspace of $\operatorname{Mat}(m\times n,\mathbb{F}_{q})$ such that rows of each matrix are in $C$. Then by [17, Proposition 5.27], there is an isomorphism from $C{[q^{m}]}$ into $S_{m,n}(C)$. For a detailed discussion, we refer the reader to [17, 23].
Theorem 3.5.
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$ and let $f\in\operatorname{Harm}_{d}(n)$. Then for any positive integer $m$, we have
$A_{i,f}^{q^{m}}(C)=\sum_{r=0}^{k}[m,r]_{q}A_{i,f}^{(r)}(C)\,.$
More specifically, we can write
$Z_{C,f}(x,y;q^{m})=\sum_{r=0}^{k}[m,r]_{q}Z_{C,f}^{(r)}(x,y)\,.$
Proof.
Let $\mathbf{u}\in C{[q^{m}]}$ and $M\in S_{m,n}(C)$ be the corresponding $m\times n$ matrix under a given isomorphism $C[q^{m}]\to S_{m,n}(C)$. Let $D_{M}\subseteq C$ be the subcode generated by the rows of $M$. Then $\operatorname{supp}(\mathbf{u})=\operatorname{Supp}(D_{M})$, and hence $\operatorname{wt}(\mathbf{u})=\operatorname{wt}(D_{M})$ (see [17, Lemma 5.4]). Therefore,
$A_{i,f}^{q^{m}}(C)$ $=\sum_{\mathbf{u}\in C{[q^{m}]},\operatorname{wt}(\mathbf{u})=i}\widetilde{f}(\operatorname{supp}(\mathbf{u}))$
$=\sum_{M\in S_{m,n}(C),\operatorname{wt}(D_{M})=i}\widetilde{f}(\operatorname{Supp}(D_{M}))$
$=\sum_{r=0}^{k}\,\sum_{\begin{subarray}{c}M\in S_{m,n}(C),\\ \mathrm{rk}({M})=r\end{subarray}}\,\sum_{\begin{subarray}{c}D_{M}\subseteq C,\\ \operatorname{wt}(D_{M})=i\end{subarray}}\widetilde{f}(\operatorname{Supp}(D_{M}))$
$=\sum_{r=0}^{k}\,\sum_{\begin{subarray}{c}M\in S_{m,n}(C),\\ \mathrm{rk}({M})=r\end{subarray}}\,\sum_{\begin{subarray}{c}D_{M}\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D_{M})=i\end{subarray}}\widetilde{f}(\operatorname{Supp}(D_{M}))$
$=\sum_{r=0}^{k}[m,r]_{q}A_{i,f}^{(r)}(C)$
This completes the proof. ∎
We can also express the harmonic higher weight enumerators in terms of the harmonic extended weight enumerators. The following result gives such correspondence.
Theorem 3.6.
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$ and let $f\in\operatorname{Harm}_{d}(n)$. Then for any integer $r$ such that $0\leq r\leq k$, we have the following relation:
$Z_{C,f}^{(r)}(x,y)=\dfrac{1}{[r]_{q}}\sum_{j=0}^{r}{\left[\!\begin{array}[]{c}r\\ j\end{array}\!\right]}_{q}(-1)^{r-j}q^{\binom{r-j}{2}}Z_{C,f}(x,y;q^{j}).$
Proof.
By Theorem 3.2, Lemmas 2.2 and 2.1 and Proposition 3.3,
$Z_{C,f}^{(r)}(x,y)$ $=(-1)^{d}\sum_{t=d}^{n-d}B_{t,f}^{(r)}(C)(x-y)^{t-d}y^{n-t-d}$
$=(-1)^{d}\sum_{t=d}^{n-d}\,\sum_{X\in E_{t}}\widetilde{f}(X)B_{X}^{(r)}(C)(x-y)^{t-d}y^{n-t-d}$
$=(-1)^{d}\sum_{t=d}^{n-d}\,\sum_{X\in E_{t}}\widetilde{f}(X){\left[\!\begin{array}[]{c}\ell(X)\\ r\end{array}\!\right]}_{q}(x-y)^{t-d}y^{n-t-d}$
$=(-1)^{d}\sum_{t=d}^{n-d}\,\sum_{X\in E_{t}}\widetilde{f}(X)\frac{[\ell(X),r]_{q}}{[r]_{q}}(x-y)^{t-d}y^{n-t-d}$
$=(-1)^{d}\frac{1}{[r]_{q}}\sum_{t=d}^{n-d}\,\sum_{X\in E_{t}}\,\sum_{j=0}^{r}{\left[\!\begin{array}[]{c}r\\ j\end{array}\!\right]}_{q}(-1)^{r-j}q^{\binom{r-j}{2}}(q^{\ell(X)})^{j}\widetilde{f}(X)$
$\hskip 150.0pt(x-y)^{t-d}y^{n-t-d}$
$=(-1)^{d}\frac{1}{[r]_{q}}\sum_{j=0}^{r}{\left[\!\begin{array}[]{c}r\\ j\end{array}\!\right]}_{q}(-1)^{r-j}q^{\binom{r-j}{2}}\sum_{t=d}^{n-d}\,\sum_{X\in E_{t}}(q^{j})^{\ell(X)}\widetilde{f}(X)$
$\hskip 150.0pt(x-y)^{t-d}y^{n-t-d}$
$=\frac{1}{[r]_{q}}\sum_{j=0}^{r}{\left[\!\begin{array}[]{c}r\\ j\end{array}\!\right]}_{q}(-1)^{r-j}q^{\binom{r-j}{2}}Z_{C,f}(x,y;q^{j}).$
4. A Greene-type theorem
In this section, we present a Greene-type theorem that relates harmonic Tutte polynomials and harmonic higher weight enumerators (resp. harmonic extended weight enumerators). As an application of this relation, we give a proof of the MacWilliams-type identities given in Theorem 3.1 (resp. Theorem 3.3). For the definitions and basic properties of Tutte polynomials of a matroid associated with the (discrete) harmonic function, we refer the reader to [11].
Definition 4.1.
Let $M=(E,\mathcal{I})$ be a matroid with rank function $\rho$ and let $f\in\operatorname{Harm}_{d}(n)$. Then the harmonic Tutte polynomial of $M$ associated to $f$ is defined as follows:
$T(M,f;x,y):={\sum_{X\subseteq E}}\widetilde{f}(X)(x-1)^{\rho(E)-\rho(X)}(y-1)^{|X|-\rho(X)}.$
By Remark 3.4, the following relation between the harmonic Tutte polynomial of a matroid and that of its dual is immediately clear:
$T(M^{\ast},f;x,y)=(-1)^{d}T(M,f;y,x)\,.$
Now we have the following Greene-type identity for harmonic extended weight enumerators.
Theorem 4.1.
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$ and let $f\in\operatorname{Harm}_{d}(n)$. Then
$Z_{C,f}(x,y;q^{m})=(-1)^{d}(x-y)^{k-d}y^{n-k-d}T\left(M_{C},f;\dfrac{x+(q^{m}-1)y}{x-y},\dfrac{x}{y}\right).$
Proof.
The proof follows very similar argument to those in [11, Theorem 4.2]. We therefore leave the details to the reader. ∎
Proof of Theorem 3.3.
Let $C$ be an $[n,k]$ code, and $M_{C}$ be its matroid. Let $k^{\perp}$ be the dimension of $C^{\perp}$. Then by Theorem 4.1,
$(-1)^{d}$ $q^{-(k-d)m}Z_{C,f}({x+(q^{m}-1)y},{x-y};q^{m})$
$=\frac{(-1)^{2d}}{q^{(k-d)m}}(q^{m}y)^{k-d}(x-y)^{n-k-d}T\left(M_{C},f;\dfrac{x}{y},\dfrac{x+(q^{m}-1)y}{x-y}\right)$
$=(x-y)^{n-k-d}y^{k-d}T\left(M_{C},f;\dfrac{x}{y},\dfrac{x+(q^{m}-1)y}{x-y}\right)$
$=(-1)^{d}(x-y)^{(n-k)-d}y^{n-(n-k)-d}T\left(M_{C^{\perp}},f;\dfrac{x+(q^{m}-1)y}{x-y},\dfrac{x}{y}\right)$
$=(-1)^{d}(x-y)^{k^{\perp}-d}y^{n-k^{\perp}-d}T\left(M_{C^{\perp}},f;\dfrac{x+(q^{m}-1)y}{x-y},\dfrac{x}{y}\right)$
$=Z_{C^{\perp},f}(x,y)\,.\qed$
Proof of Theorem 3.1.
By Theorems 3.6 and 3.3, we can write
$Z_{C^{\perp},f}^{(r)}(x,y)$ $=\dfrac{1}{[r]_{q}}\sum_{j=0}^{r}{\left[\!\begin{array}[]{c}r\\ j\end{array}\!\right]}_{q}(-1)^{r-j}q^{\binom{r-j}{2}}Z_{C^{\perp},f}(x,y;q^{j})$
$=(-1)^{d}\dfrac{1}{[r]_{q}}\sum_{j=0}^{r}{\left[\!\begin{array}[]{c}r\\ j\end{array}\!\right]}_{q}(-1)^{r-j}q^{\binom{r-j}{2}}\frac{(q^{j})^{d}}{q^{jk}}$
$\hskip 100.0ptZ_{C,f}\left({x+(q^{j}-1)y},{x-y};q^{j}\right)$
$=\dfrac{1}{[r]_{q}}\sum_{j=0}^{r}{\left[\!\begin{array}[]{c}r\\ j\end{array}\!\right]}_{q}(-1)^{r+d-j}q^{\binom{r-j}{2}-j(k-d)}$
$\hskip 100.0ptZ_{C,f}\left({x+(q^{j}-1)y},{x-y};q^{j}\right)$
$=\dfrac{1}{[r]_{q}}\sum_{j=0}^{r}{\left[\!\begin{array}[]{c}r\\ j\end{array}\!\right]}_{q}(-1)^{r+d-j}q^{\binom{r-j}{2}-j(k-d)}\sum_{\ell=0}^{j}[j,\ell]_{q}$
$\hskip 100.0ptZ_{C,f}^{(\ell)}\left({x+(q^{j}-1)y},{x-y}\right)$
$=\dfrac{1}{[r]_{q}}\sum_{j=0}^{r}\,\sum_{\ell=0}^{j}{\left[\!\begin{array}[]{c}r\\ j\end{array}\!\right]}_{q}(-1)^{r+d-j}q^{\binom{r-j}{2}-j(k-d)}[j,\ell]_{q}$
$\hskip 100.0ptZ_{C,f}^{(\ell)}\left({x+(q^{j}-1)y},{x-y}\right).$
The proof is completed by the following easily-proven identity:
$\dfrac{1}{[r]_{q}}{\left[\!\begin{array}[]{c}r\\ j\end{array}\!\right]}_{q}[j,\ell]_{q}=\frac{1}{q^{j(r-j)}[r-j]_{q}q^{\ell(j-\ell)}[j-\ell]_{q}}\,.\qed$
Example 4.1.
Let $C$ be a $[5,2]$ code over $\mathbb{F}_{2}$ with generator matrix:
$\begin{pmatrix}1&1&1&0&0\\ 0&0&0&1&1\end{pmatrix}.$
Then the dual code $C^{\perp}$ has the generator matrix:
$\begin{pmatrix}1&0&1&0&0\\ 0&1&1&0&0\\ 0&0&0&1&1\end{pmatrix}.$
We consider $f\in\operatorname{Harm}_{1}(5)$, where $f=a\{1\}+b\{2\}+c\{3\}+d\{4\}-(a+b+c+d)\{5\}$. By direct calculation, we have
$W_{C,f}^{(0)}$ $=0,$
$W_{C,f}^{(1)}$ $=(a+b+c)(xy)^{1}(xy^{2}-x^{2}y),$
$W_{C,f}^{(2)}$ $=0.$
Here $Z_{C,f}^{(0)}=Z_{C,f}^{(2)}=0$ and $Z_{C,f}^{(1)}=(a+b+c)(xy^{2}-x^{2}y)$. It is evident that $C^{\perp}$ has seven subcodes of dimension $2$. The explicit list of subcodes is given in [6]. Now by Theorem 3.1, we have
$Z_{C^{\perp},f}^{(2)}(x,y)$ $=\sum_{j=0}^{2}\sum_{\ell=0}^{j}(-1)^{3-j}\frac{2^{{2-j\choose 2}-j(2-j)-\ell(j-\ell)-j}}{[2-j]_{2}[j-\ell]_{2}}Z_{C,f}^{(\ell)}\left({x+(2^{j}-1)y},{x-y}\right)$
$=(a+b+c)(xy^{2}-y^{3}).$
Hence $W_{C^{\perp},f}^{(2)}(x,y)=(a+b+c)(xy)(xy^{2}-y^{3})$.
The following remarks concerning harmonic generalization of the MacWilliams identity for $r$-th higher weight enumerators stated in Theorem 3.1 might be helpful:
Remark 4.1.
Using the binomial series expansion for $i=d,\ldots,n-d$, we obtain the identity
$(x+(q^{j}-1)y)^{n-d-i}(x-y)^{i-d}$ $=\sum_{p=d}^{n-d}K_{p}(i-d;n-2d,q^{j})x^{n-d-p}y^{p-d},$
where $K_{p}(i-d;n-2d,q^{j})$ is known as the Krawtchouk polynomial. For any positive integer $n>2d$, it is defined as follows (see [23]):
$K_{p}(i-d;n-2d,q^{j}):=\sum_{m=0}^{p}(-1)^{m}(q^{j}-1)^{p-m}\binom{i-d}{m}\binom{n-d-i}{p-m}.$
Remark 4.2.
Since $A_{i,f}^{(0)}(C^{\perp})=0$, we can write $A_{i,f}^{(r)}(C^{\perp})$ explicitly for $i=d,\ldots,n-d$ and $r\neq 0$ as
$A_{i,f}^{(r)}(C^{\perp})=(-1)^{d}\sum_{\ell=1}^{r}\,$ $\sum_{j=\ell}^{r}(-1)^{r-j}\frac{q^{{r-j\choose 2}-j(r-j)-\ell(j-\ell)-j(k-d)}}{[r-j]_{q}[j-\ell]_{q}}$
$\sum_{p=d}^{n-d}K_{p}(i-d;n-2d,q^{j})A_{p,f}^{(\ell)}(C).$
Theorem 4.2.
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$ and let $f\in\operatorname{Harm}_{d}(n)$. Then we have the following Greene-type relation:
$Z_{C,f}^{(r)}(x,y)=\frac{1}{[r]_{q}}\sum_{j=0}^{r}{\left[\!\begin{array}[]{c}r\\ j\end{array}\!\right]}_{q}(-1)^{r+d-j}q^{\binom{r-j}{2}}$ $(x-y)^{k-d}y^{n-k-d}$
$T\left(M_{C},f;\dfrac{x+(q^{j}-1)y}{x-y},\dfrac{x}{y}\right).$
Proof.
Apply Theorems 3.6 and 4.1. ∎
5. Applications to subcode designs
Britz and Shiromoto [8] gave a generalization of the Assmus-Mattson Theorem [1] by introducing designs from subcode supports of an $\mathbb{F}_{q}$-linear code; see Theorem 5.2 below. In this section, we provide a new proof of this generalization by using the harmonic higher weight enumerators of an $\mathbb{F}_{q}$-linear code. Our method of proof is inspired by Bachoc [2] who first used the concept of harmonic functions to prove the Assmus-Mattson Theorem for $t$-designs.
A $t$-$(n,k,\lambda)$ design is a collection $\mathcal{B}$ of $k$-subsets (called blocks) of a set $E$ of $n$ elements, such that any $t$-subset of $E$ is contained in exactly $\lambda$ blocks. A design is called simple if the blocks are distinct; otherwise, the design is said to have repeated blocks.
The following theorem gives a remarkable characterization of $t$-designs in terms of the harmonic spaces.
Theorem 5.1 ([14]).
Let $i,t$ be integers such that $0\leq t\leq i\leq n$. A subset $\mathcal{B}\subseteq E_{i}$ is a $t$-design if and only if $\sum_{B\in\mathcal{B}}(\sum_{A\in E_{d},A\subseteq B}f(A))=0$ for all $f\in\operatorname{Harm}_{d}(n)$, $1\leq d\leq t$.
Example 5.1.
Let $C$ be a $[6,3]$ code over $\mathbb{F}_{2}$ with generator matrix
$\begin{pmatrix}1&1&0&0&0&0\\ 0&0&1&1&0&0\\ 0&0&0&0&1&1\end{pmatrix}.$
There are seven linear subcodes of $C$ with dimension $2$. Three of them have weight $4$; the set of supports of these three is
$\mathcal{S}_{2,4}=\{\{1,2,3,4\},\{1,2,5,6\},\{3,4,5,6\}\}\,.$
Note that $\mathcal{S}_{2,4}(C)$ is a $1$-design. Moreover, for $f\in\operatorname{Harm}_{1}(6)$, where $f=a\{1\}+b\{2\}+c\{3\}+d\{4\}+e\{5\}-(a+b+c+d+e)\{6\}$, we get $A_{4,f}^{(2)}(C)=0$.
Now we have the following useful lemma.
Lemma 5.1.
Let $C$ be an $[n,k]$ code over $\mathbb{F}_{q}$. Let $r,i$ be the positive integers such that $r\leq k$ and $i\geq d_{r}$. Then the set $\mathcal{S}_{r,i}(C)$ forms a $t$-design if and only if $A_{i,f}^{(r)}(C)=0$ for all $f\in\operatorname{Harm}_{d}(n)$, $1\leq d\leq t$.
Proof.
By the definition of $\widetilde{f}$ from (2), we have
$A_{i,f}^{(r)}(C)$ $=\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D)=i\end{subarray}}\widetilde{f}(\operatorname{Supp}(D))$
$=\sum_{\begin{subarray}{c}D\in\mathcal{D}_{r}(C),\\ \operatorname{wt}(D)=i\end{subarray}}\sum_{\begin{subarray}{c}Z\in E_{d},\\ Z\subseteq\operatorname{Supp}(D)\end{subarray}}f(Z)$
$=\sum_{Y\in\{\operatorname{Supp}(D)\mid D\in\mathcal{D}_{r}(C),\operatorname{wt}(D)=i\}}\sum_{\begin{subarray}{c}Z\in E_{d},Z\subseteq Y\end{subarray}}f(Z)$
$=\sum_{Y\in\mathcal{S}_{r,i}(C)}\sum_{\begin{subarray}{c}Z\in E_{d},Z\subseteq Y\end{subarray}}f(Z).$
Hence, the proof follows by Theorem 5.1. ∎
Below is the subcode generalization of the Assmus-Mattson Theorem.
Theorem 5.2 ([8]).
Let $C$ be an $[n,k,d_{min}]$ code over $\mathbb{F}_{q}$. Let $r,t\geq 1$ be integers with $r\leq k$ and $t\leq d_{min}$. Define
$\mathcal{L}_{r,t}:=\bigg{\{}i\in\{d_{r}^{\perp},\ldots,n-t\}\;\bigg{|}\;{\sum_{\ell=0}^{r}[r]_{\ell}\,A_{i}^{(\ell)}(C^{\perp})}\neq 0\bigg{\}},$
where $d_{r}^{\perp}$ is the $r$-th generalized Hamming weight of the dual code $C^{\perp}$. Suppose that $|\mathcal{L}_{\mu,t}|\leq d_{\mu}-t$ for all $\mu=1,\ldots,r$. Then for all $i\geq d_{r}$ and $j=d_{r}^{\perp},\ldots,n-t$, $\mathcal{S}_{r,i}(C)$ and $\mathcal{S}_{r,j}(C^{\perp})$ each form a (not necessarily simple) $t$-design.
Proof.
Let $f\in\operatorname{Harm}_{d}(n)$ for $1\leq d\leq t$. By Lemma 5.1, it is sufficient to show that $A_{i,f}^{(r)}(C)=0$ (resp. $A_{i,f}^{(r)}(C^{\perp})=0$) for all $i=d,\ldots,n-d$. Applying Theorem 3.1 and Remark 4.2 to the dual code $C^{\perp}$ gives the identity
(6) $(-1)^{d}\sum_{\ell=1}^{r}M_{\ell}^{r}\underline{A}^{(\ell)}(C^{\perp};f)=\underline{A}^{(r)}(C;f),$
where $\underline{A}^{(\ell)}(C^{\perp};f):=(A_{d,f}^{(\ell)}(C^{\perp}),\ldots,A_{n-d,f}^{(\ell)}(C^{\perp}))$ (resp. $\underline{A}^{(r)}(C;f):=(A_{d,f}^{(r)}(C),\ldots,A_{n-d,f}^{(r)}(C))$) and where $M_{\ell}^{r}$ is the square matrix of order $n-2d+1$ defined for all $i=d,\ldots,n-d$ and $p=0,\ldots,n-2d$ by
$(M_{\ell}^{r})_{p,i-d}:=\sum_{j=\ell}^{r}(-1)^{r-j}\frac{q^{{r-j\choose 2}-j(r-j)-\ell(j-\ell)-j(k^{\perp}-d)}}{[r-j]_{q}[j-\ell]_{q}}K_{p}(i-d;n-2d,q^{j}),$
where $k^{\perp}$ denotes the dimension of $C^{\perp}$ and $K_{p}(i-d;n-2d,q^{j})$ is the Krawtchouk polynomial given in Remark 4.1. Since $d_{r}$ is the $r$-th generalized Hamming weight of $C$,
(7) $A_{d,f}^{(r)}(C)=A_{d+1,f}^{(r)}(C)=\cdots=A_{d_{r}-1,f}^{(r)}(C)=0\,.$
For $r=1$, this leads to $d_{1}-d$ independent equations in the terms $A_{i,f}^{(1)}(C^{\perp})$ for $d\leq i\leq n-d$. Since $|\mathcal{L}_{1,d}|\leq d_{1}-d$ and $1\leq d\leq t\leq d_{\min}$, there are at most $d_{1}-d$ unknowns and so the only solution is trivial. Hence, $A_{i,f}^{(1)}(C^{\perp})=0$ for all $i=d,\ldots,n-d$.
Similarly, the $n-2d+1$ equations $A_{i,f}^{(1)}(C^{\perp})=0$, $d\leq i\leq n-d$, each translate into an equation in the terms $A_{i,f}^{(1)}(C)$, $d\leq i\leq n-d$, using equation (6) applied to the dual code $C^{\perp}$. Since these equations are independent, the only solution is trivial; that is, $A_{i,f}^{(1)}(C)=0$ for all $i=d,\ldots,n-d$.
Now let $r=m+1$ for any positive integer $m$ such that $11\leq m\leq k-$ 1. Then by induction, we have from (7) that there are $d_{m+1}-d$ independent equations in the terms $A_{i,f}^{(m+1)}(C^{\perp})$ for $d\leq i\leq n-d$ as
$A_{d,f}^{(\ell)}(C^{\perp})=A_{d+1,f}^{(\ell)}(C^{\perp})=\cdots=A_{n-1,f}^{(\ell)}(C^{\perp})=0$
for all $1\leq\ell\leq m$. Hence, by a similar argument to that above, $A_{i,f}^{(m+1)}(C^{\perp})=0$ (resp. $A_{i,f}^{(m+1)}(C)=0$) for all $i=d,\ldots,n-d$. This completes the proof. ∎
Acknowledgements
This work was supported by JSPS KAKENHI (22K03277) and SUST Research Centre (PS/2023/1/22). The authors would also like to thank the anonymous reviewers for their beneficial comments on an earlier version of the manuscript.
Data availability statement
The data that support the findings of this study are available from the corresponding author.